# S2malloc: Statistically Secure Allocator for Use-After-Free Protection And More

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Software security engineering
1. Introduction
Heap-related vulnerabilities are serious and common threats that can be leveraged to launch attacks resulting in arbitrary code execution or information leakage. Heap overflow, double and invalid free, and use-after-free (UAF) are among the most common types of these vulnerabilities. According to the Common Vulnerabilities and Exposures (CVEs) report of 2022, they were ranked 19th, 11th, and 7th respectively, in terms of bug prevalence (National Vulnerability Database, [n. d.]).
@@ 1. Introduction
UAF occurs when a previously freed memory block is used to store data. It receives special attention in secure allocator design due to its prevalence and the powerful exploitation primitives (e.g. arbitrary read/write) it enables. Chromium has reported that more than one-third of their security issues are related to UAF, more prevalent than other types of memory errors combined (The Chromium Projects, [n. d.]). While complete mitigation of UAF is theoretically possible by virtually never re-using a previously freed memory block, UAF-mitigating allocators incur substantial overheads (e.g., about 40% for MarkUs (Ainsworth and Jones, 2020)), and their complexity leaves them vulnerable to new attacks (Wickman et al., 2021) (albeit preventable).
On the other hand, entropy-based heap allocators aim to provide comprehensive protection against most, if not all, common heap vulnerabilities with simpler designs but may fail with a small probability. Specifically on UAF mitigation, entropy-based memory allocators typically use delayed free-lists (Silvestro et al., 2018; Liu et al., 2019) to prevent the same memory block from being immediately re-allocated after being freed. Attackers now face a moving-target even when they manage to obtain a dangling pointer as they have less confidence in knowing when this pointer becomes valid again and/or which object it might point to. While achieving relatively low overhead on time usage, especially compared with UAF-mitigating allocators (see § 2 for details), existing entropy-based allocators still face the challenge of entropy-loss due to heap spraying, information leak, and silent failures on (potentially repeated) trials.
@@
Interested readers can contact the authors for the code.
2. Background and Related Work
In this section, we give a succinct overview of common heap vulnerabilities and the most recent advancements in two defense lines: moving-target defense against generic heap exploitations and complete mitigation against use-after-free (UAF)—the current most frequently encountered type of memory error (The Chromium Projects, [n. d.]).
2.1. Heap vulnerabilities
@@
2.2. Entropy-based allocators
Entropy-based heap allocators strive to provide protections against almost all of the above-mentioned heap exploitations by minimizing the attacker’s success rate. Ideally, the success rate should be low enough to deter attackers from even trying to attack the system. However, practical implementations face limitations in terms of memory and computation resources, as demonstrated in Guarder (Silvestro et al., 2018).
@@ 2.2. Entropy-based allocators
BIBOP.BIBOPBIBOP. State-of-the-art entropy-based allocators typically build on the Big Bag of Pages (BIBOP) (Hanson, 1980) memory management mechanism with various security enhancements. BIBOP-style allocators classify allocation sizes as classes. For each size class, one or several continuous pages are treated as a bag, and all allocations of the same size class will be assigned to a corresponding bag. Each bag is split into several slots preemptively and each allocated object will take one of them. The status of each slot is monitored by a bitmap and can be used to defend against double or invalid frees. Bag metadata is stored separately to avoid metadata-based attacks (Shellphish, [n. d.]), and UAF only occurs within the same size class.
@@
DieHarder (Novark and Berger, 2010) is one of the earliest entropy-based secure allocators that adopts the BIBOP-style memory management and provides statistical protections against heap exploitations. Despite incorporating all the security features mentioned above, DieHarder has several issues compared to more recent works. These include unstable allocation entropy, predictable guard pages, and a significant impact on the execution time of the hardened program.
Guarder (Silvestro et al., 2018) offers multiple improvements over DieHarder, such as using a linked-list-based free block management algorithm instead of bitmap traversal, providing stable allocation entropy, and offering in-bag tunable random guard pages. Guarder significantly reduces the run-time overhead to less than 3%, making it suitable for production systems.
SlimGuard (Liu et al., 2019) further extends Guarder by reducing its memory overhead. It divides size classes into 16-byte increments instead of powers of two. However, the current implementation of SlimGuard has two security compromises: it 1) allocates blocks sequentially in the free-list, violating the claim of random allocation. 2) reuses freed blocks to store metadata, possibly due to an implementation flaw, violating the argued metadata separation practice.
2.3. UAF-mitigating allocators
Conventional heap allocators prioritize defragmentation and memory saving and hence, tend to eagerly re-allocate an object to a recently freed memory block of a similar size. An attacker can easily abuse this allocation strategy to predict future allocations. In other words, this defragmentation strategy favors the attackers and makes UAF exploitation (§ 2.1) easier. The eager reallocation strategy is based on implicitly trusting that the free() operation issued by the developer is an accurate indication of the end of the lifetime of the memory object. In reality, however, people make mistakes and a free() call might be issued when there are still dangling references.
UAF-mitigating allocators address this issue by simply not treating free() as a signal to safely free the allocation. Instead, the allocator only frees a memory block when it is absolutely safe to do so. While recent works may have arguably different criteria on when it is absolutely safe to free a memory block, in theory, this style of defense could eliminate all UAF exploitations, even in the presence of heap spraying and/or repeated trials. As a result, these allocators can be considered as a complete defense against UAF.
@@ 2.3. UAF-mitigating allocators
Recent works.Recent worksRecent works. Following are representative examples of allocators adopting the complete mitigation line of thinking:
Pointer Tracking and Nullification: MarkUs (Ainsworth and Jones, 2020) adopts a Garbage Collection (GC) style to mitigate UAF by reusing memory slots only after no references remain. Freed memory is initially quarantined in a free list and will be released after they are no longer accessible, confirmed through periodically checking the registers and memory for accessible live objects. MineSweeep (Erdős et al., 2022) further extends MarkUs by using a simple linear sweep of memory to detect dangling pointers. DangNull (Lee et al., 2015) takes a similar approach but instead of counting the number of dangling pointers, it changes their pointed memory to NULL. Cornucopia (Filardo et al., 2020) uses a sweeping technique that nullifies dangling pointers with only introducing less than 2% overheads. However, Cornucopia requires the support of the CHERI (Watson et al., 2015) architecture, which is not available in commercial CPUs. FFmalloc (Wickman et al., 2021) provides an extreme solution to mitigate UAF by never reusing the virtual memory address and unmapping all freed memory. As there is no memory re-use at all, UAF can never happen. Oscar (Dang et al., 2017) modifies the page permission on deallocation thus stopping subsequent access via dangling pointers.
Context-sensitive mitigations: Cling (Akritidis, 2010) only reuses freed blocks and allocates them to objects of the same type, which is inferred from the call stack address. PUMM (Yagemann et al., 2023) analyzes the program call flow graph (CFG) and releases a freed block to be reused after the previous task (e.g., a request for a network server) is finished. These allocators are also considered complete UAF mitigation within their corresponding contexts.
@@
Privileged allocators.Privileged allocatorsPrivileged allocators. Besides userspace solutions, UAF-mitigating allocators requiring privilged access have also been proposed. For example, DangZero (Gorter et al., 2022) detects UAF by assigning a physical memory address with multiple virtual memory address aliases. The current alias will be marked invalid after being freed and a new alias will be used to access it if this physical memory is reallocated.
DangZero requires a privileged backend to get access to the direct page table. PTAuth (Farkhani et al., 2020) uses Pointer Authentication Code (PAC) (Arm Limited, [n. d.]), to generate signatures for legitimate objects. Signatures are checked upon dereference and are invalidated upon deallocation. PTAuth requires a kernel patch to store PAC keys.
2.4. Widely deployed secure allocators
The industry has adopted security designs into their allocators. Scudo (LLVM Project, 1024) is a solution from LLVM (LLVM Project, [n. d.]) and supports random allocation and delayed free-list. Hardened_malloc (GrapheneOS, 2024) is the default allocator of GrapheneOS (GrapheneOS, [n. d.]), a privacy and security-focused Android-based OS. While applying all aforementioned designs in entropy-based allocators and providing UAF-write protections, it suffers from at least reduced randomness and increased overheads for larger blocks, making the allocation predictable and limiting its use scenarios.
3. Motivation
@@ 3.1. Adversary model
We do, however, assume that the underlying OS kernel and hardware are trusted and an attacker cannot utilize a data leakage channel, such as /proc/\$pid/maps, to discover the location of the heap allocator’s metadata. The attacker cannot compromise the random number generator nor can they take control of the heap allocator. Exploiting bugs of the allocator itself is out of scope. These assumptions are similar to that of other entropy-based allocators (Silvestro et al., 2018; Liu et al., 2019).
Additionally, we allow attackers to use any existing heap feng-shui (Sotirov, 2007) technique (e.g., heap spray) to prepare or manipulate the layout of heap to facilitate UAF-exploits. And attackers can retry an exploit as long as previous attempts fail silently. These assumptions make our adversary model stronger than those assumed in entropy-based allocators and on-par with the adversary models in UAF-mitigating allocators (Ainsworth and Jones, 2020; Wickman et al., 2021).
3.2. Challenge 1: entropy loss
@@
Figure 1 is an example, abstracted from mRuby issue 4001 (Daniel Teuchert, Cornelius Aschermann, Tommaso Frassetto, Tigist Abera, 2018), a UAF vulnerability in the Ruby compiler. The function mrbioinitializecopy is called when opening a file. It first frees the existing data pointer of the copy object (DATAPTR(copy)) (line 7) and allocates new memory to it (line 9 and 11). If an invalid argument is passed, calling iogetopenfptr would throw an exception (line 10), making DATAPTR(copy) a dangling pointer to the freed object.
@@
[0   ] mrbioinitializecopy(mrbstate *mrb, mrbvalue copy) mrbvalue orig; struct mrbio *fptrcopy, *fptrorig;
fptrcopy = (struct mrbio *)DATAPTR(copy); if (fptrcopy != NULL) mrbfree(mrb, fptrcopy); fptrcopy = (struct mrbio *)mrbioalloc(mrb); fptrorig = iogetopenfptr(mrb, orig); DATAPTR(copy) = fptrcopy;
Figure 1. Example UAF attack based on mRuby issue 4001 (Daniel Teuchert, Cornelius Aschermann, Tommaso Frassetto, Tigist Abera, 2018)
In the above scenario, random (i.e., non-sequential) allocation or delayed free-lists available in existing entropy-based allocators (Novark and Berger, 2010; Silvestro et al., 2018; Liu et al., 2019) merely increases the attack difficulty: as long as the attacker can wait, the memory chunk referred to by the dangling pointer will eventually be re-allocated, allowing the UAF exploit to proceed after some delay.
@@
3.3. Challenge 2: information leak
Existing entropy-based allocators (Silvestro et al., 2018; Liu et al., 2019) create a huge memory pool for each block size range, resulting in the leakage of block size via their address, possibly revealing the victim software’s internal state to the attacker. For example, each BIBOP bag is assigned an 8 Gigabytes virtual memory pool in SlimGuard (Liu et al., 2019). Any objects belonging to this bag will be allocated from this pool. As a result, obtaining a known sized block will be sufficient enough to infer the size of any blocks sharing the upper 32-bit address. Further, block addresses in Guarder (Silvestro et al., 2018) are aligned by their size. Block size would possibly be inferred just based on its address. For example, a block with an address value ending with 0x10 is of size 1 to 16 bytes, and a block with an address value ending with 0x300 is highly likely of size 129 to 256 bytes.
@@ 4.1. Architectural overview
4.2. Randomized in-slot offset (RIO)
Figure 2. Overview of S2malloc with an example of free and malloc. A, B, and C show three S2malloc segments, stored in segregated memory. 1, 2, and 3 show how an allocated bag slot is freed and then allocated.
@@ 4.2. Randomized in-slot offset (RIO)
We define $k$ to be the RIO entropy. For each bag with blocks of $b$ bytes, $e=b/k$ bytes are not used to take data, and each block can take at most $b-e$ bytes of data, guarantee minimal in-slot offset entropy. We refer to these extra bytes as entropy bytes. Suppose this block is malloc-ed with an object of $s$ bytes ($s<b-e$). Before this block is allocated, $p\in(0,b-s)$ is computed to decide the starting byte of the data object. $p$ is 16-byte aligned following the minimum default alignment of GNU C implementation (The GNU Project, [n. d.]), and to be compatible with special data structures, such as atomic objects that need to be stored to align with the registers and cache lines. The offset of each block will be stored separately in an offset table. The minimal entropy $e$ increases as $b$ becomes larger to avoid introducing high memory overhead for small objects and to provide stronger protections for large objects, observing the fact that larger objects have more complicated structures and are more likely to be targeted.
We adopt the Permuted Congruential Generator (PCG) (O’Neill, 2014) algorithm following the design of  (Liu et al., 2019) to generate all random numbers. With negligible execution time – much faster than existing random number generators, such as LCG and Unix XorShift, PCG generates hard-to-predict numbers. While it does not provide cryptographic security, the only known attack towards PCG requires three consecutive PCG outputs to recover the seed (Bouillaguet et al., 2020). However, in S2malloc, none of the generated numbers is accessible to the user and can only be obtained by sweeping the memory. As a result, PCG is sufficient enough to provide the required level of entropy.
4.3. Random bag layout (RBL)
@@
S2malloc obfuscates the virtual memory allocation and stops linking block sizes to their addresses. Instead of allocating a dedicated virtual memory pool for each bag, (as shown in prior works (Silvestro et al., 2018; Liu et al., 2019)), we create a single virtual memory address pool for all bags. We further divide each bag into sub-bags each containing 256 slots. Each bag creates new sub-bags upon need, and a newly created sub-bag would request corresponding memory from the pool. We use a bump pointer to track the available memory in the pool and linearly allocate pool memory to sub-bags.
@@
Secure
Allocators
Guard
Pages
Rand.
Alloc.
Segre.
MD
Heap
Canary
Pointer
Inval.
UAF
Mitigation
UAF
Detection
Memory
Overhead
Runtime
Overhead
DieHarder (Novark and Berger, 2010) Y Y Y N N One-shot attack only N 21.3% 2.1%
Guarder (Silvestro et al., 2018) Y Y Y Y N One-shot attack only N 58.1% 2.4%
SlimGuard (Liu et al., 2019) Y N N Y N One-shot attack only N 22.5% 4.4%
S2malloc Y Y Y Y N Repeated/spray attacks High prob. 26.8% 2.8%
MarkUs (Ainsworth and Jones, 2020) N N N N Y Repeated/spray attacks N ${}^{\ast}$ ${}^{\ast}$
FFmalloc (Wickman et al., 2021) N N N N Y Repeated/spray attacks Y ${}^{\ast}$ ${}^{\ast}$
Table 1. Overview of existing secure memory allocators and S2malloc to illustrate how S2malloc fills the gap (MD: metadata). Memory and run-time overheads are measured by running the PARSEC benchmark (Bienia et al., 2008). Note that overheads of MarkUs and FFmalloc (numbers marked with $\ast$) are reported in (Wickman et al., 2021) instead of measured by us. The details for getting other overhead numbers are presented in § 6. With that said, the performance numbers shown here are for a qualitative illustration on the scale of overhead only. For quantitative comparisons, please refer to details in § 6.
4.4. Hardening heap canaries
Canary is a small data block put after the allocated memory to detect overflow, initially introduced in StackGuard (Cowan et al., 1998) to protect the stack. Canary has now been adopted to protect the heap (Nikiforakis et al., 2013). At the time a memory slot is being allocated, the canary will be set to a specific value. This value will be checked at the time this slot is freed, and memory overflow will be detected if the canary value changes.
However, in previous designs, this value is set to be either globally identical (Silvestro et al., 2018) or is binded with slots (Liu et al., 2019), and could be trivially broken by a knowledgeable attacker. We follow the design of previous works to use the secure MAC of the memory address as the canary (Liljestrand et al., 2019; Mashtizadeh et al., 2015). Specifically, we take the CMAC-AES-128 encrypted block address as the canary implemented using AES-NI (Intel Corporation, [n. d.]) (on x86 CPUs) or Neon (ARM Developer, [n. d.]) (on AARCH CPUs) to keep the canary confidential and compact. Even if the attacker learns a canary value, they can only use it to break the current object or any further objects allocated to this slot with the same RIO.

### v2

@@ Abstract
Keywords:
Secure Memory Allocator Use-After-Free.
1 Introduction
Heap-related vulnerabilities are serious and common threats that can be leveraged to launch attacks resulting in arbitrary code execution or information leakage. Heap overflow, double and invalid free, and use-after-free (UAF) are among the most common types of these vulnerabilities. According to the Common Vulnerabilities and Exposures (CVEs) report of 2022, they were ranked 19th, 11th, and 7th respectively, in terms of bug prevalence [27].
@@ 1 Introduction
UAF occurs when a previously freed memory block is used to store data. It receives special attention in secure allocator design due to its prevalence and the powerful exploitation primitives (e.g. arbitrary read/write) it enables. Chromium has reported that more than one-third of their security issues are related to UAF, more prevalent than other types of memory errors combined [38]. While complete mitigation of UAF is theoretically possible by virtually never re-using a previously freed memory block, UAF-mitigating allocators incur substantial overheads (e.g., about 40% for MarkUs [3]), and their complexity leaves them vulnerable to new attacks [42] (albeit preventable).
On the other hand, entropy-based heap allocators aim to provide comprehensive protection against most, if not all, common heap vulnerabilities with simpler designs but may fail with a small probability. Specifically on UAF mitigation, entropy-based memory allocators typically use delayed free-lists [33, 24] to prevent the same memory block from being immediately re-allocated after being freed. Attackers now face a moving-target even when they manage to obtain a dangling pointer as they have less confidence in knowing when this pointer becomes valid again and/or which object it might point to. While achieving relatively low overhead on time usage, especially compared with UAF-mitigating allocators (see § 2 for details), existing entropy-based allocators still face the challenge of entropy-loss due to heap spraying, information leak, and silent failures on (potentially repeated) trials.
@@
The software artifact is open-sourced.
2 Background and Related Work
In this section, we give a succinct overview of common heap vulnerabilities and the most recent advancements in two defense lines: moving-target defense against generic heap exploitations and complete mitigation against use-after-free (UAF)—the current most frequently encountered type of memory error [38].
2.1 Heap vulnerabilities
@@
2.2 Entropy-based allocators
Entropy-based heap allocators strive to provide protections against almost all of the above-mentioned heap exploitations by minimizing the attacker’s success rate. Ideally, the success rate should be low enough to deter attackers from even trying to attack the system. However, practical implementations face limitations in terms of memory and computation resources, as demonstrated in Guarder [33].
@@ 2.2 Entropy-based allocators
BIBOP.BIBOPBIBOP. State-of-the-art entropy-based allocators typically build on the Big Bag of Pages (BIBOP) [17] memory management mechanism with various security enhancements. BIBOP-style allocators classify allocation sizes as classes. For each size class, one or several continuous pages are treated as a bag, and all allocations of the same size class will be assigned to a corresponding bag. Each bag is split into several slots preemptively and each allocated object will take one of them. The status of each slot is monitored by a bitmap and can be used to defend against double or invalid frees. Bag metadata is stored separately to avoid metadata-based attacks [32], and UAF only occurs within the same size class.
@@
DieHarder [29] is one of the earliest entropy-based secure allocators that adopts the BIBOP-style memory management and provides statistical protections against heap exploitations. Despite incorporating all the security features mentioned above, DieHarder has several issues compared to more recent works. These include unstable allocation entropy, predictable guard pages, and a significant impact on the execution time of the hardened program.
Guarder [33] offers multiple improvements over DieHarder, such as using a linked-list-based free block management algorithm instead of bitmap traversal, providing stable allocation entropy, and offering in-bag tunable random guard pages. Guarder significantly reduces the run-time overhead to less than 3%, making it suitable for production systems.
SlimGuard [24] further extends Guarder by reducing its memory overhead. It divides size classes into 16-byte increments instead of powers of two. However, the current implementation of SlimGuard has two security compromises: it 1) allocates blocks sequentially in the free-list, violating the claim of random allocation. 2) reuses freed blocks to store metadata, possibly due to an implementation flaw, violating the argued metadata separation practice.
2.3 UAF-mitigating allocators
UAF-mitigating allocators specialize in UAF-protection only and can be broadly categorized into the following types:
  • A
    Pointer validation on dereference [40, 14, 43]: i.e., checks whether a pointer is valid and safe for a read/write operation upon dereference.
  • B
    Pointer invalidation on free [16, 10, 31, 15, 42]: i.e., whenever free is called on a pointer, it revokes its capability to access the associated virtual address.
  • C
    Memory sweeping [20, 3, 13]: i.e., checks all pointers stored in memory and either actively removes all dangling pointers or reuses a freed memory chunk with the assurance that no dangling pointer to this freed chunk exists.
  • D
    Context-based allocation [4, 19, 12]: i.e., permits the re-allocation of freed memory chunks only to objects allocated in the same “context.”
Not all UAF-mitigating allocators are drop-in replacements of the system memory allocator, as some of them require recompilation of the protected programs [4, 19, 41], special hardware [15], or kernel modifications [16].
@@ 2.3 UAF-mitigating allocators
Security guarantee.Security guaranteeSecurity guarantee. UAF-mitigating allocators in categories A, B, and C can eliminate all UAF exploits (hence complete mitigation), although some of them might incur large overheads or require customized kernel or hardware. Context-based allocators (category D), first proposed in [12] as a heap-safety property, typically run faster but offer incomplete protection. For example, Cling [4] defines the allocation context based on the two innermost return addresses on the call stack when malloc is invoked. Suppose two memory allocations occur in the same context; in this case, the object allocated the second time may occupy the same memory chunk allocated the first time (if the first object is freed). TypeAfterType [19] defines the allocation context based on object types (e.g., the type int passed in malloc(sizeof(int)), and objects may be reallocated on freed memory chunks originating from the same allocation context. In summary, UAF is still possible in context-based allocators.
However, unlike entropy-based allocators (including S2malloc), which also mitigate UAF imperfectly, the protection from context-based allocators is deterministic. Specifically, if the dangling pointer and the target object are allocated in different contexts, there is zero chance of success in corrupting the target object, regardless of the attackers’ strategies. On the other hand, context-based allocators provide no protection if the dangling pointer and the target object fall in the same context. Entropy-based allocators provide probabilistic protection in both cases. We will discuss the implications via CVEs in § 5.5.
@@
Porting for UAF-write defense only.Porting for UAF-write defense onlyPorting for UAF-write defense only. Allocators that instrument runtime checks for pointer validity (i.e., category A) can opt to trade protections of UAF-read attacks for performance by checking the pointer validity on write accesses only. In fact, as reported in [34], only 5% to 25% of memory accesses are write accesses in the SPEC 2017 benchmark, indicating a potential reduction in overhead for category-A allocators. On the other hand, for allocators in categories B, C, and D, splitting UAF-read and UAF-write protection is inherently hard as their designs do not differentiate between read and write accesses.
3 Motivation
@@ 3.1 Adversary model
We do, however, assume that the underlying kernel and hardware are trusted and an attacker cannot utilize a data leakage channel, like /proc/\$pid/maps, to discover the location of the heap allocator’s metadata. The attacker cannot compromise the random number generator nor can they take control of the heap allocator. Exploiting bugs of the allocator itself is out of scope. These assumptions are similar to that of other entropy-based allocators [33, 24].
Additionally, we allow attackers to use any existing heap feng-shui [35] technique (e.g., heap spray) to prepare or manipulate the layout of heap to facilitate UAF-exploits. And attackers can retry an exploit as long as previous attempts fail silently. These assumptions make our adversary model stronger than those assumed in entropy-based allocators and on-par with the adversary models in UAF-mitigating allocators [3, 42].
3.2 Challenge 1: entropy loss
@@
Figure 1 is an example, abstracted from mRuby issue 4001 [11], a UAF vulnerability in the Ruby compiler. The function mrbioinitializecopy is called when opening a file. It first frees the existing data pointer of the copy object (DATAPTR(copy)) (line 7) and allocates new memory to it (line 9 and 11). If an invalid argument is passed, calling iogetopenfptr would throw an exception (line 10), making DATAPTR(copy) a dangling pointer to the freed object.
@@
1mrb_io_initialize_copy(mrb_state *mrb, mrb_value copy) {
2  mrb_value orig;
3  struct mrb_io *fptr_copy, *fptr_orig;
4
5  fptr_copy = (struct mrb_io *)DATA_PTR(copy);
6  if (fptr_copy != NULL) {
7    mrb_free(mrb, fptr_copy);
8  }
9  fptr_copy = (struct mrb_io *)mrb_io_alloc(mrb);
10  fptr_orig = io_get_open_fptr(mrb, orig);
11  DATA_PTR(copy) = fptr_copy;
12}
Figure 1: Example UAF attack based on mRuby issue 4001 [11]
In the above scenario, random (i.e., non-sequential) allocation or delayed free-lists available in existing entropy-based allocators [29, 33, 24] merely increases the attack difficulty: as long as the attacker can wait, the memory chunk referred to by the dangling pointer will eventually be re-allocated, allowing the UAF exploit to proceed after some delay.
@@ 11  DATA_PTR(copy) = fptr_copy;
3.3 Challenge 2: information leak
Existing entropy-based allocators [33, 24] create a huge memory pool for each block size range, resulting in the leakage of block size via their address, possibly revealing the victim software’s internal state to the attacker. For example, each BIBOP bag is assigned an 8 Gigabytes virtual memory pool in SlimGuard [24]. Any objects belonging to this bag will be allocated from this pool. As a result, obtaining a known sized block will be sufficient enough to infer the size of any blocks sharing the upper 32-bit address. Further, block addresses in Guarder [33] are aligned by their size. Block size would possibly be inferred just based on its address. For example, a block with an address value ending with 0x10 is of size 1 to 16 bytes, and a block with an address value ending with 0x300 is highly likely of size 129 to 256 bytes.
@@ 4.1 Architectural overview
4.2 Randomized in-slot offset (RIO)
Figure 2: Overview of S2malloc with an example of free and malloc. A, B, and C show three S2malloc segments, stored in segregated memory. 1, 2, and 3 show how an allocated bag slot is freed and then allocated.
@@ 4.2 Randomized in-slot offset (RIO)
We define $k$ to be the RIO entropy. For each bag with blocks of $b$ bytes, $e=b/k$ bytes are not used to take data, and each block can take at most $b-e$ bytes of data, guarantee minimal in-slot offset entropy. We refer to these extra bytes as entropy bytes. Suppose this block is malloc-ed with an object of $s$ bytes ($s<b-e$). Before this block is allocated, $p\in(0,b-s)$ is computed to decide the starting byte of the data object. $p$ is 16-byte aligned following the minimum default alignment of GNU C implementation [39], and to be compatible with special data structures, such as atomic objects that need to be stored to align with the registers and cache lines. The offset of each block will be stored separately in an offset table. The minimal entropy $e$ increases as $b$ becomes larger to avoid introducing high memory overhead for small objects and to provide stronger protections for large objects, observing the fact that larger objects have more complicated structures and are more likely to be targeted.
We adopt the Permuted Congruential Generator (PCG) [30] algorithm following the design of  [24] to generate all random numbers. With negligible execution time – much faster than existing random number generators, such as LCG and Unix XorShift, PCG generates hard-to-predict numbers. While it does not provide cryptographic security, the only known attack towards PCG requires three consecutive PCG outputs to recover the seed [7]. However, in S2malloc, none of the generated numbers is accessible to the user and can only be obtained by sweeping the memory. As a result, PCG is sufficient enough to provide the required level of entropy.
4.3 Random bag layout (RBL)
@@
S2malloc obfuscates the virtual memory allocation and stops linking block sizes to their addresses. Instead of allocating a dedicated virtual memory pool for each bag, (as shown in prior works [33, 24]), we create a single virtual memory address pool for all bags. We further divide each bag into sub-bags each containing 256 slots. Each bag creates new sub-bags upon need, and a newly created sub-bag would request corresponding memory from the pool. We use a bump pointer to track the available memory in the pool and linearly allocate pool memory to sub-bags.
@@
Secure Guard Rand. Segre. Heap Ptr UAF UAF Overheads
Allocators Pages Alloc. MD Canary Inval. Mitigate Detect Memory Runtime
DieHarder Y Y Y N N N 21.3% 2.1%
Guarde Y Y Y Y N N 58.1% 2.4%
SlimGuard Y N N Y N N 22.5% 4.4%
S2malloc Y Y Y Y N High prob. 26.8% 2.8%
MarkUs N N N N Y N 13.0% 42.9%
FFmalloc N N N N Y Y 50.5% 33.1%
Table 1: Overview of existing secure memory allocators and S2malloc to illustrate how S2malloc fills the gap (MD: metadata, ◐: One-shot attack only, ●: Repeated/spray attacks). Overheads are measured on PARSEC [6], detailed in § 6. Overheads of MarkUs and FFmalloc (numbers marked with $\ast$) are reported in [42] instead of measured by us.
4.4 Hardening heap canaries
Canary is a small data block put after the allocated memory to detect overflow, initially introduced in StackGuard [8] to protect the stack. Canary has now been adopted to protect the heap [28]. At the time a memory slot is being allocated, the canary will be set to a specific value. This value will be checked at the time this slot is freed, and memory overflow will be detected if the canary value changes.
However, in previous designs, this value is set to be either globally identical [33] or is binded with slots [24], and could be trivially broken by a knowledgeable attacker. We follow the design of previous works to use the secure MAC of the memory address as the canary [21, 26]. Specifically, we take the CMAC-AES-128 encrypted block address as the canary implemented using AES-NI [18] (on x86 CPUs) or Neon [5] (on AARCH CPUs) to keep the canary confidential and compact. Even if the attacker learns a canary value, they can only use it to break the current object or any further objects allocated to this slot with the same RIO.